OpenStudy (anonymous):

Solve for x: −2(x + 3) = −2(x + 1) − 4

OpenStudy (anonymous):

I was left with 0=1, sound right? heres what i did...

OpenStudy (nurali):

−2(x + 3) = −2(x + 1) − 4 -2x-6=-2x-2-4 -2x-6=-2x-6 -2x+2x-6+6=0 0=0 All are real number.

OpenStudy (anonymous):